# Resource Allocation for Twin Maintenance and Computing Task Processing in Digital Twin Vehicular Edge Computing Network

### v1

@@ III-C Optimization Problem
Due to the fact that vehicle $i$ simultaneously performs digital twin maintenance and offloading of computing tasks at time slot $t$, VEC server will allocate computing resources for both, namely $f_{i}^{dt}$ and $f_{i}^{tk}$, respectively. By using different computing resource allocation strategies, different information synchronization delays and task computation delays can be obtained, which in turn affects the maintenance delay $t_{i,B}^{dt}$ of digital twins and the processing delay $t_{i,B}^{tk}$ of computing tasks. According to [62], we define a satisfaction function $Q$ to demonstrate the satisfaction level of digital twin maintenance delay and computational task processing delay for vehicle $i$ under a certain resource allocation strategy $\omega_{i}$, respectively. Therefore, the satisfaction function $Q_{i}^{dt}(\omega_{i})$ for the maintenance delay of vehicle $i$ digital twin and the satisfaction function $Q_{i}^{tk}(\omega_{i})$ for the processing delay of the computational task for vehicle $i$ can be respectively represented as
$Q_{i}^{dt}(\omega_{i})=1-\ln(1+\frac{t_{i}^{dt}}{T_{i}^{dt}}),$ (9)
$Q_{i}^{tk}(\omega_{i})=1-\ln(1+\frac{t_{i}^{tk}}{T_{i}^{tk}}),$ (10)
Through (9) and (10), we can evaluate the resource utility $U_{i}(\omega_{i})$ obtained by vehicle $i$ under a resource allocation strategy $\omega_{i}$
$U_{i}(\omega_{i})=\rho Q_{i}^{dt}(\omega_{i})+(1-\rho)Q_{i}^{tk}(\omega_{i}),$ (11)
where $\rho$ is the weighting factor with $10<\rho<$ 1 between twin maintenance and computational task processing, used to measure the importance of the two tasks.
@@
${\rm{}}s.t.\quad{\rm{}}t_{i}^{dt}\leq T_{i}^{dt},\$ (12b)
$\qquad{\rm{}}t_{i}^{tk}\leq T_{i}^{tk},\$ (12c)
$\qquad\qquad\quad\sum_{i=1}^{N}f_{i}^{dt}+f_{i}^{tk}\leq F,\$ (12d)
@@ IV Proposed solution
Each vehicle in the scene will simultaneously perform digital twin maintenance and offloading of computing tasks to the server. The computing resources requested by a vehicle to the server to complete these two tasks will also affect the resources obtained by other vehicles, thereby affecting the final utility. In addition, the computing resources requested by each vehicle can be either integer or real variables. Obviously, our optimization problem is a mixed integer nonlinear programming problem, which is NP-hard and cannot be completed in polynomial time using traditional methods. Therefore, we transformed the problem into a multi-agent MDP and proposed a MARL-CSTC algorithm to solve it, while analyzing the complexity of the algorithm.
@@ IV-A2 State Space $\mathcal{S}$
The state $s_{n}(t)$ of agent $n$ at moment $t$ can be described by updating information, computing tasks, vehicle position, speed, and channel gain, which can be represented as $s_{n}(t)=\{\Gamma_{n}(t),\zeta_{n}(t),PO_{n}(t),v_{n}\}$. The state of the entire system $s(t)$ can be expressed as $s(t)=\{s_{n}(t)\}_{N}$.
@@ IV-A3 Action Space $\mathcal{A}$
The behavior $a_{n}(t)$ taken by the agent $n$ includes computing resources for the maintenance of the digital twin of agent $n$, $f_{n}^{dt}$, as well as resources for the processing of computing tasks of agent $n$, $f_{n}^{tk}$. Therefore, the behavior $a_{n}(t)$ taken by agent $n$ can be expressed as $a_{n}(t)=\{f_{n}^{dt},f_{n}^{tk}\}$. The action space $a(t)$ for all agents can be represented as $a(t)=\{a_{n}(t)\}_{N}$.
@@ IV-A5 Reward Function $\mathcal{R}$
$r_{n}(t)=\rho Q_{n}^{dt}(\omega_{n})+(1-\rho)Q_{n}^{tk}(\omega_{n}),$ (13)
@@
Figure 3: Structure of MARL-CSTC
@@ IV-B Algorithm Design
As shown in Figure 3, for the structure of MARL-CSTC algorithm, actor-critic framework is employed, where the actor is used to generate the actions that the agent needs to perform, while the critic is responsible for guiding the actor network to generate better actions. Here, the actor consists of two parts, where the estimation actor network is responsible for training, and the target actor network is responsible for executing the action. Similarly, critic also includes two parts, namely the estimation critic network and the target critic network, both of which are used to evaluate the actions of agents. For this reason, the actor network adopts a policy based deep neural network, while the critic uses a value based deep neural network. In addition, considering the dynamic changes in the environment, we adopt a centralized training and distributed execution strategy, where the critic is centrally trained by the server and the actor is executed by each vehicle in a distributed manner.
@@
For the estimation actor network of agent $n$, its input is the current state $s_{n}(t)$, including the synchronization information required for twin maintenance $\Gamma_{n}(t)$, computational tasks generated by vehicles $\theta_{n}(t)$, vehicle position $PO_{n}(t)$, and vehicle speed $v_{n}$. The current state is input into the fully connected layer, processed by three hidden layers, and output as the probability of all possible actions. After passing through the softmax layer, the total probability of all actions is set to 1. Finally, agent $n$ selects one of them as the final resource allocation strategy to be executed.
@@
Algorithm 1 Training Stage of MARL-CSTC Algorithm.
0:  $N,D_{n}^{dt},C_{n}^{dt},T_{n}^{dt},D_{n}^{tk},C_{n}^{tk},T_{n}^{tk},v_{n},g_{n,B}^{t}$ for $n=1,...,N$;
@@ IV-C Algorithm Training
The centralized training of the MARL-CSTC algorithm is implemented on the server, as summarized in Algorithm 1. Specifically, since the server manages the critic network of each agent, the status and actions of all agents can be obtained, meaning that the information of all agents is observable. In the centralized training phase, the server first obtains the states and actions of all agents. By utilizing this information, the server can train the estimation critical network for each agent, thereby achieving the goal of maximizing the Q value. For a single agent $n$, its Q-value $Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})$ can be updated through the Bellman equation[63], which is given by $Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})=R_{n}(t)+\gamma_{n}\max Q_{n}^{{}^{\prime}}(\mathcal{S}^{{}^{\prime}},\mathcal{A}^{{}^{\prime}};\theta^{{}^{\prime}}_{Q_{n}})$. The time difference error can be calculated as $\delta=Q_{n}^{{}^{\prime}}(\mathcal{S}^{{}^{\prime}},\mathcal{A}^{{}^{\prime}};\theta^{{}^{\prime}}_{Q_{n}})-Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})$. Therefore, the loss function can be expressed as $L(\theta_{Q_{n}})=\mathbb{E}(\delta^{2})$. To minimize the loss function $L(\theta_{Q_{n}})$, we update the parameters $\theta_{Q_{n}}$ by using a stochastic gradient descent algorithm, which is represented as $\nabla_{\theta_{Q_{n}}}L(\theta_{Q_{n}})=\mathbb{E}(2\delta\nabla_{\theta_{Q_{n}}}Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}}))$.
@@
Algorithm 2 Executing Stage of MARL-CSTC Algorithm.
@@ IV-D Algorithm Execution
After the centralized training is completed, distributed actions need to be executed on each vehicle based on local observations. Specifically, agent $n$ downloads the results of centralized training and inputs them into the actor network. Then, agent $n$ observes and obtains the local state $s_{n}(t)$, and executes action $a_{n}(t)$ based on policy $\pi_{n}$ to obtain a reward $r_{n}(t)$. In addition, due to a lack of experience in the early stages of execution, agent $n$ randomly selects actions to explore. When experience is sufficient, it will execute actions to maximize rewards. Therefore, in order to achieve a balance between exploration and exploitation, we adopt the $\epsilon$-greedy algorithm, as summarized in Algorithm 2.
@@ IV-E Algorithm Complexity Analysis
Inspired by [37], the computational complexity of algorithms is mainly related to the structure and parameters of neural networks. Due to the fact that both the actor network and the critic network of the MARL-CSTC algorithm use DNN, the calculation of complexity needs to be based on the analysis of DNN. We let $Ln$ represent the number of layers in DNN, and let $G_{l}$ represent the number of neurons in the $l$-th layer. Therefore, we can obtain that the computational complexity of the actor network and the critic network is $\mathcal{O}(X_{a})=\mathcal{O}(X_{c})=\mathcal{O}(\sum_{l=1}^{Ln}G_{l}G_{l+1})$.
@@ V Simulation results
In this section, to evaluate the performance of the MARL-CSTC algorithm on the scenario considered in this paper, we conducted simulation experiments and compared it with other algorithms to analyze its superiority.
@@ V-A Experimental Setup
The simulation experiment in this article was implemented using Python 3.11.0 and Pytorch. In terms of environmental settings, the experimental parameters are listed in Table $\mathrm{I}$. The number of lanes $J$ is set to 3, the bandwidth $W$ is set to 150MHZ, the noise powers $\sigma^{2}$ are set to $10^{-11}mW$ and the information required for twin maintenance and the calculation task size generated by the vehicle itself are set to uniformly sample between [1024,1536] bytes. The CPU frequency required to process these two units of Btye is set to $C_{i}^{dt}=C_{i}^{tk}=0.25$MHZ/Byte, with a maximum delay tolerance of 0.5s. We set the channel correlation coefficient $\kappa$ to 0.2, the path loss index $\beta$ to 3, and set the weight $\rho$ to 0.5, which indicates that twin maintenance and vehicular computing task processing are equally important. In terms of DNN, for actor networks, the number of neurons in the first two hidden layers of the three fully connected layers is 300 and 100, respectively, and the number of neurons in the third hidden layer is the action dimension. For the critic network, the number of neurons in the three hidden layers is 300, 100 and 1, respectively. During the training phase, in terms of learning rate, the actor network has a learning rate of $lr_{a}=2\times 10^{-4}$, the critic network has a learning rate of $lr_{c}=10^{-3}$, and a discount factor $\gamma_{n}=0.95$.
@@
$W$ 150MHZ
$\sigma^{2}$ $10^{-11}mW$
$D_{i}^{dt},D_{i}^{tk}$ [1024,1536]Bytes
@@
Figure 4: Average rewards under different algorithms
@@ V-B Algorithm Comparison
We compared three algorithms, namely Soft Actor-Critic (SAC) algorithm, Proximal Policy Optimization (PPO) algorithm, and Random algorithm.
  • Both SAC algorithm and PPO algorithm can be regarded as single agent algorithms.At each epoch, only one agent updates its actions based on locally obtained information and trained DQN, while the actions of other agents remain unchanged. Single DQN shared among all agents.
  • Random algorithm is the process of first dividing the total computing resources by the number of agents at each decision time t, and then randomly allocating computing resources for twin maintenance and processing computing tasks to the corresponding agents within each average component.
@@
(b) The ratio of resource conversion to utility
Figure 5: Comparison of resource utilization and the ratio of resource conversion to utility under different algorithms
(a) The impact of the number of vehicles on resource utility:$p_{i}^{max}$=200mW
(b) The impact of maximum transmission power on resource utility:$N$=5
Figure 6: Comparison of resource utility under different algorithms

### v2

@@ III-D Optimization Problem
Due to the fact that vehicle $i$ simultaneously performs digital twin maintenance and offloading of computing tasks at time slot $t$, vehicle $i$ will request computing resources for both, namely $f_{i}^{\text{dt}}$ and $f_{i}^{\text{tk}}$, respectively. By using different resource request strategies, different information synchronization delays and task computing delays can be obtained, which in turn affects the maintenance delay $t_{i,B}^{\text{dt}}$ of digital twins and the processing delay $t_{i,B}^{\text{tk}}$ of computing tasks. According to [44], we define a satisfaction function $Q$ to demonstrate the satisfaction level of digital twin maintenance delay and computing task processing delay for vehicle $i$ under a certain resource request strategy $\omega_{i}$, respectively. Therefore, the satisfaction function $Q_{i}^{\text{dt}}(\omega_{i})$ for the maintenance delay of vehicle $i$ digital twin and the satisfaction function $Q_{i}^{\text{tk}}(\omega_{i})$ for the processing delay of the computational task for vehicle $i$ can be respectively represented as
$Q_{i}^{\text{dt}}(\omega_{i})=1-\ln(1+\frac{t_{i}^{\text{dt}}}{T_{i}^{\text{dt}}}),$ (11)
$Q_{i}^{\text{tk}}(\omega_{i})=1-\ln(1+\frac{t_{i}^{\text{tk}}}{T_{i}^{\text{tk}}}),$ (12)
The satisfaction function reflects the level of satisfaction with the current resource allocation strategy, as the twin maintenance latency and task processing latency are higher, the value of the satisfaction function is lower, which indicates that the current strategy is poor. Therefore, through (9) and (10), we can evaluate the resource utility $U_{i}(\omega_{i})$ obtained by vehicle $i$ under a resource allocation strategy $\omega_{i}$
$U_{i}(\omega_{i})=\rho Q_{i}^{\text{dt}}(\omega_{i})+(1-\rho)Q_{i}^{\text{tk}}(\omega_{i}),$ (13)
where $\rho$ is the weighting factor with $10<\rho<$ 1 between twin maintenance and computational task processing, used to measure the importance of the two tasks. The weight parameters are set to be fixed here because the vehicle environment considered in this paper is a smooth state in an ideal environment. However, it is important to note that dynamic weights are more suitable to cope with emergency mission situations.
@@
${\rm{}}s.t.\quad{\rm{}}t_{i}^{\text{dt}}\leq T_{i}^{\text{dt}},\$ (14b)
$\qquad{\rm{}}t_{i}^{\text{tk}}\leq T_{i}^{\text{tk}},\$ (14c)
$\qquad\qquad\quad\sum_{i=1}^{N}(f_{i}^{\text{dt}}+f_{i}^{\text{tk}})\leq F,\$ (14d)
@@ IV Proposed solution
Each vehicle in the scene will simultaneously perform digital twin maintenance and offloading of computing tasks to the server. The computing resources requested by a vehicle to complete these two tasks will also affect the resources obtained by other vehicles, thereby affecting the final utility. We transformed the problem into a multi-agent MDP and proposed a TMTPRCS algorithm to solve it, while analyzing the complexity of the algorithm. The algorithm refers to MADDPG and is partially optimized based on it. Specifically, the traditional MADDPG algorithm relies on injecting Gaussian noise for exploration, which can be inefficient and lacks dynamic adaptation. In contrast, TMTPRCS employs an adaptive $\epsilon$-greedy algorithm that dynamically adjusts throughout the training process, promoting exploration during the early stages and gradually shifting towards exploitation as training progresses. Unlike conventional greedy algorithms, TMTPRCS ensures that effective actions and corresponding knowledge are not lost during the exploration phase.
@@ IV-A2 State Space $\mathcal{S}$
The state $s_{n}(t)$ of agent $n$ at moment $t$ can be described by updating information, computing tasks, vehicle position, speed, and channel gain, which can be represented as $s_{n}(t)=\{\Gamma_{n}(t),\zeta_{n}(t),PO_{n}(t),v_{n},g_{n,B}^{t}\}$. The state of the entire system $s(t)$ can be expressed as $s(t)=\{s_{n}(t)\}_{N}$.
@@ IV-A3 Action Space $\mathcal{A}$
The behavior $a_{n}(t)$ taken by the agent $n$ includes request for resources for twin maintenance $f_{n}^{\text{dt}}$ and computing task processing $f_{n}^{\text{tk}}$. Therefore, the behavior $a_{n}(t)$ can be expressed as $a_{n}(t)=\{f_{n}^{\text{dt}},f_{n}^{\text{tk}}\}$. The action space $a(t)$ for all agents can be represented as $a(t)=\{a_{n}(t)\}_{N}$.
@@ IV-A5 Reward Function $\mathcal{R}$
$r_{n}(t)=\rho Q_{n}^{\text{dt}}(\omega_{n})+(1-\rho)Q_{n}^{\text{tk}}(\omega_{n}),$ (15)
@@
By maximizing the long-term cumulative reward of each agent, i.e., $\max R_{n}(t)=\sum_{t_{0}}^{t}{\gamma_{n}r_{n}(t)}$, we can then obtain the optimal resource allocation strategy.
Figure 3: Structure of TMTPRCS
@@ IV-B Algorithm Design
As shown in Figure 3, for the structure of TMTPRCS algorithm, actor-critic framework is employed, where the actor is used to generate the actions that the agent needs to perform, while the critic is responsible for guiding the actor network to generate better actions. Here, the actor consists of two parts, where the estimation actor network is responsible for training, and the target actor network is responsible for executing the action. Similarly, critic also includes two parts, namely the estimation critic network and the target critic network, both of which are used to evaluate the actions of agents. For this reason, the actor network adopts a policy based deep neural network, while the critic uses a value based deep neural network. In addition, considering the dynamic changes in the environment, we adopt a centralized training and distributed execution strategy, where the critic is centrally trained by the server and the actor is executed by each vehicle in a distributed manner. It is noted that centralized training is done in an offline manner and distributed execution is online.
@@
For the estimation actor network of agent $n$, its input is the current state $s_{n}(t)$, including the synchronization information required for twin maintenance $\Gamma_{n}(t)$, computational tasks generated by vehicles $\theta_{n}(t)$, vehicle position $PO_{n}(t)$, vehicle speed $v_{n}$, and channel gain $g_{n,B}^{t}$. The current state is input into the fully connected layer, processed by three hidden layers, and output as the probability of all possible actions. After passing through the softmax layer, the total probability of all actions is set to 1. Finally, agent $n$ selects one of them as the final resource allocation strategy to be executed.
@@
Algorithm 1 Training Stage of TMTPRCS Algorithm.
0:  $N,D_{n}^{\text{dt}},C_{n}^{\text{dt}},T_{n}^{\text{dt}},D_{n}^{\text{tk}},C_{n}^{\text{tk}},T_{n}^{\text{tk}},v_{n},g_{n,B}^{t}$ for $n=1,...,N$;
@@ IV-C Algorithm Training
The centralized training of the TMTPRCS algorithm is implemented on the server, as summarized in Algorithm 1. Specifically, since the server manages the critic network of each agent, the status and actions of all agents can be obtained, meaning that the information of all agents is observable. In the centralized training phase, the server first obtains the states and actions of all agents. By utilizing this information, the server can train the estimation critical network for each agent, thereby achieving the goal of maximizing the Q value. For a single agent $n$, its Q-value $Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})$ can be updated through the Bellman equation[55], which is given by $Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})=R_{n}(t)+\gamma_{n}\max Q_{n}^{{}^{\prime}}(\mathcal{S}^{{}^{\prime}},\mathcal{A}^{{}^{\prime}};\theta^{{}^{\prime}}_{Q_{n}})$. The time difference error can be calculated as $\delta=Q_{n}^{{}^{\prime}}(\mathcal{S}^{{}^{\prime}},\mathcal{A}^{{}^{\prime}};\theta^{{}^{\prime}}_{Q_{n}})-Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})$. Therefore, the loss function can be expressed as $L(\theta_{Q_{n}})=\mathbb{E}(\delta^{2})$. To minimize the loss function $L(\theta_{Q_{n}})$, we update the parameters $\theta_{Q_{n}}$ by using a stochastic gradient descent algorithm, which is represented as $\nabla_{\theta_{Q_{n}}}L(\theta_{Q_{n}})=\mathbb{E}(2\delta\nabla_{\theta_{Q_{n}}}Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}}))$.
@@
Algorithm 2 Executing Stage of TMTPRCS Algorithm.
@@ IV-D Algorithm Execution
After the centralized training is completed, distributed actions need to be executed on each vehicle based on local observations. Specifically, agent $n$ downloads the results of centralized training and inputs them into the actor network. Then, agent $n$ observes and obtains the local state $s_{n}(t)$, and executes action $a_{n}(t)$ based on policy $\pi_{n}$ to obtain a reward $r_{n}(t)$. In addition, due to a lack of experience in the early stages of execution, agent $n$ randomly selects actions to explore. When experience is sufficient, it will execute actions to maximize rewards. Therefore, in order to achieve a balance between exploration and exploitation, we adopt the $\epsilon$-greedy algorithm. In addition, we have optimized $\epsilon$ so that it can be used to maintain a balance between exploration and exploitation as training progresses. It can be expressed as $\epsilon=(1-\lambda)^{K}\epsilon_{0}$, where $\lambda$ is rate of descent, $\epsilon_{0}$ is initial value of $\epsilon$. The detailed process is as summarized in Algorithm 2.
@@ IV-E Algorithm Complexity Analysis
Similar to the analysis process in [30], the computational complexity of algorithms is mainly related to the structure and parameters of neural networks. Due to the fact that both the actor network and the critic network of the TMTPRCS algorithm use DNN, the calculation of complexity needs to be based on the analysis of DNN. We let $Ln$ represent the number of layers in DNN, and let $G_{l}$ represent the number of neurons in the $l$-th layer. Therefore, we can obtain that the computational complexity of the actor network and the critic network is $\mathcal{O}(X_{a})=\mathcal{O}(X_{c})=\mathcal{O}(\sum_{l=1}^{Ln}G_{l}G_{l+1})$.
@@ V Simulation results
In this section, to evaluate the performance of the TMTPRCS algorithm on the scenario considered in this paper, we conducted simulation experiments and compared it with other algorithms to analyze its superiority.
@@ V-A Experimental Setup
The simulation experiment in this article was implemented using Python 3.11.0 and Pytorch. In terms of environmental settings, the experimental parameters are listed in Table $\mathrm{I}$. The number of lanes $J$ is set to 3. The distance $L_{0}$ between first lane and BS and the width $\omega_{0}$ of each lane are 8m and 3.75m, respectively. The duration $\tau$ of each time slot is set to 100ms. The bandwidth $W$ between vehicle $i$ and BS is set to 20Mhz, the transmission power of vehicle $i$ is set to 23dBm and the noise power $\sigma^{2}$ is -110dBm. The carrier frequency $f_{c}$ is 2Ghz. The speed $v_{i}$ of vehicle $i$ randomly sample between [10,15]m/s. We assume that the information required for twin maintenance is the vehicle control task, whose size $D_{i}^{\text{dt}}$ and time constraints $T_{i}^{\text{dt}}$ are set to uniformly sample between [10,200] bytes and 100ms, respectively. The size $D_{i}^{\text{tk}}$ of the computing task generated by the vehicle is set to uniformly sample between [1000, 1500]bytes, and its time limit $T_{i}^{\text{tk}}$ is 500ms. The cpu frequency $C_{i}^{dt}/C_{i}^{tk}$ required per unit byte is set to 0.25Mhz/byte. In terms of DNN, for actor networks, the number of neurons in the first two hidden layers of the three fully connected layers is 300 and 100, respectively, and the number of neurons in the third hidden layer is the action dimension. For the critic network, the number of neurons in the three hidden layers is 300, 100 and 1, respectively. The learning rates for the actor network and critic network are $lr_{a}=10^{-4}$ and $lr_{c}=10^{-3}$, respectively. The discount factor $\gamma_{n}$ and parameter update rate $\eta$ are set to 0.95 and 0.01. It is noted that simulation parameters are fixed to ensure fair baseline comparisons. We acknowledge that the sensitivity of TMTPRCS to variations in environmental and algorithmic parameters remains an open question, and plan to conduct a detailed sensitivity analysis in future work to evaluate robustness under diverse network settings.
@@
$L_{0}$ 8m
$\omega_{0}$ 3.75m
$\tau$ 100ms
$W$ 20Mhz
$p_{i}$ 23dBm
$\sigma^{2}$110dBm
$f_{c}$ 2Ghz
$v_{i}$ [10,15]m/s
$D_{i}^{\text{dt}}$ [10,200]bytes
$D_{i}^{\text{tk}}$ [1000, 1500]bytes
$T_{i}^{\text{dt}}$ 100ms
$T_{i}^{\text{tk}}$ 500ms
@@
$\eta$ 0.01
Figure 4: Average rewards under different algorithms:$N=7$
@@ V-B Algorithm Comparison
We compared four algorithms, namely Multi-Agent Twin Delayed Deep Deterministic(MATD3) algorithm, Policy Gradient Soft Actor-Critic(SAC) algorithm, Proximal Policy Optimization(PPO) algorithm, and Random algorithm.
  • MATD3: A Multi-agent Algorithm for improving stability and performance by introducing dual Q networks and delayed updates of target policies.
  • SAC: An algorithm based on maximum entropy reinforcement learning aims to maximize the cumulative reward while maximizing the entropy of the strategy to encourage exploration.
  • PPO: An algorithm that ensures training stability by limiting the magnitude of policy updates.
  • Random: We make a simple modification to the Random algorithm to satisfy the constraint 14d. Specifically, Random algorithm is the process of first dividing the total computing resources by the number of agents at each decision time $t$, and then randomly allocating computing resources for twin maintenance and processing computing tasks to the corresponding agents within each average component.
@@
(b) Schedulable resources for twin maintenance
(c) Schedulable resources for tasks processing
Figure 5: Comparison of resource utilization and schedulable resources for twin maintenance/task processing under different algorithms
Figure 6: Utility under different algorithms